OpenStudy (anonymous):

find the root of 6x - 48= 0 I got x = -8 check my answer?

hartnn (hartnn):

-8 or +8 ?

hartnn (hartnn):

6x -48 = 0 6x = 48 ...

OpenStudy (anonymous):

it would be -48

OpenStudy (anonymous):

right?

hartnn (hartnn):

6x-48 =0 you would need to ADD 48 on both sides, right ?

hartnn (hartnn):

6x-48 +48 = 0 +48 6x +0 = +48 6x = 48

OpenStudy (anonymous):

if you add 48 together its 96 so you would have to subtract it from both sides

OpenStudy (anonymous):

cause you have to cancel it out

hartnn (hartnn):

see, we need to find 'x' so we need to isolate it we have -48 with 6x to first isolate 6x and get rid of -48 we do the inverse operation of subtraction(which is addition) so that -48+48 becomes 0 so we ADD 48 on both sides (now we never get 96, because we don't have 48+48)

OpenStudy (anonymous):

ohh

hartnn (hartnn):

6x - 48 + 48 = 0+48 6x = +48

OpenStudy (anonymous):

x is 6?

hartnn (hartnn):

6x = 48 to isolate x, we divide 6 on both sides 6x/6 = 48/6 x = 4/6 = 8 :)

OpenStudy (anonymous):

nvm its x=8

hartnn (hartnn):

YES

OpenStudy (anonymous):

:)